OpenStudy (anonymous):

Can someone help explain how to get the answer (sq. root)5 / 5x out of the problem:: (sq. root) 3x / (sq. root) 15x^3 ?? I'm severely lost as to how to get that answer.

OpenStudy (anonymous):

\[\frac{\sqrt{3x}}{\sqrt{15x ^{3}}} = \frac{\sqrt{3x}}{x \sqrt{3x \times 5 }} = \frac{1}{x \sqrt{5}}\]

OpenStudy (anonymous):

multiply the numerator and denominator by square root of 5 \[\frac{1}{x \sqrt{5}} = \frac{1 \times \sqrt{5}}{x \sqrt{5} \times \sqrt{5}} = \frac{\sqrt{5}}{5x}\]

OpenStudy (anonymous):

understood?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

for that one you need to get equal denominator and then cross all the denominator because they all are same and try to simplfy the upper value the numanoter

OpenStudy (anonymous):

did u get it or no